Dear Colleague, I will like to give an update on work that has been done around the experiment in preparation for the upcoming run, particular in the last couple of weeks. New trigger counters for MRS and FS have been designed, constructed including stands. Constructed at BNL, holder/stand for FS at TAMU (D.Beavis, E.Baker, K.Hagel). These detectors have been installed on platform (D.Beavis, E.Baker) The triggers counters used for pp, dA runs has been removed, and the cabling for the Bema Beam Counters restored (D.Beavis, E.Baker, B.Budick) Signal cables for the triggers counters have been made, and the cable run to FEH installed (R.Scheetz, J.Hammond). The Si-detectors have been re-configured with two rings assigned for flow measurements. The new holders were constructed at Kansas. The electronics will be air-cooled; checkout has still to be done (S.Sanders, H.Ito). The annual safety review has been conducted, and the items noted in the report are being corrected (D.Beavis, R.Scheetz). An additional Fastbus crate has been installed in FEH to accommodated signals for new detectors; software has been installed (R.Scheetz, F.Videbaek) The procedures for shift-leaders (i.e. all) are being updated, and will be placed in binders at CH and in physics office (D.Beavis). These procedures slightly different from last year will have to be read and acknowledged by all. General cleanup of IR have been made. best regards Flemming -- Flemming Videbaek Brookhaven National Laboratory Physics Department tlf 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Fri Nov 28 16:32:02 2003
This archive was generated by hypermail 2.1.8 : Fri Nov 28 2003 - 16:32:20 EST